### v2

# Eyes Wide Unshut: Unsupervised Mistake Detection in Egocentric Video by Detecting Unpredictable Gaze

Abstract
In this paper, we address the challenge of unsupervised mistake detection in egocentric video through the analysis of gaze signals, a critical component for advancing user assistance in smart glasses. Traditional supervised methods, reliant on manually labeled mistakes, suffer from domain-dependence and scalability issues. This research introduces an unsupervised method for detecting mistakes in videos of human activities, overcoming the challenges of domain-specific requirements and the necessity for annotated data. By analyzing unusual gaze patterns that signal user disorientation during tasks, we propose a gaze completion model that forecasts eye gaze trajectories from incomplete inputs. The difference between the anticipated and observed gaze paths acts as an indicator for identifying errors. Our method is validated on the EPIC-Tent dataset, showing its superiority compared to current one-class supervised and unsupervised techniques.
Index Terms:
Gaze, Mistake, Unsupervised, Behaviour Understanding
I Introduction
Smart glasses have recently gained significant popularity, with various existing products available on the market capable of providing assistance to the user through Augmented Reality. In order to provide timely assistance, wearable devices should be able to identify moments in which the user makes mistakes or is confused and requires help [1]. If such instances are properly detected, the AI system can proactively support users by offering contextual information or suggestions on how to best carry out the task at hand [2]. Previous works tackled the problem of detecting mistakes from a fully supervised perspective, where mistake instances were labeled in egocentric video and a machine learning algorithm was trained to discriminate between video segments of correct action executions and incorrect ones [3, 2]. Such a fully supervised approach has two main downsides: 1) It is domain-dependent, hence requiring an accurate characterization of what a mistake is, depending on the context. For instance, a mistake in a kitchen scenario is different from a mistake in the assembly line; 2) It requires a sufficient number of manually labeled mistake instances, which is time consuming, requires expert knowledge and sometimes it is difficult to observe and record. Ideally, a wearable assistant should be able to infer when the behavioral patterns of the user deviate from the norm in order to determine if they need assistance in a scenario-independent setting, i.e., without making specific assumptions on how a mistake is defined and without requiring costly labels.
To overcome the limitations of previous approaches, we propose to detect mistakes in egocentric videos of human activity in an unsupervised way, by learning from unlabeled video.
Figure 1: The proposed unsupervised mistake detection method. (a) We assume as input a video with a partial gaze trajectory on the initial part of the video. (b) A gaze completion model predicts a gaze trajectory for the remaining part of the video, conditioned on the input video and the partial trajectory. (c) A mistake is detected if the predicted trajectory is significantly different from the observed one, which suggests an anomalous attention pattern.
We postulate that mistakes are characterized by abnormal behavior exhibited by the camera wearer, who may feel disoriented during the execution of the task at hand. As widely acknowledged in the literature, gaze patterns play an important role in the control of human activities [4], hence, we study how to analyze them as a means to perform mistake detection. We propose to train a model of human gaze behavior, in the form of a gaze predictor, and use it to predict likely eye gaze trajectories from video at inference time. Assuming that ground truth gaze trajectories are available as input, we frame gaze prediction as a novel “gaze completion” task in which a model takes as input a video and a partial gaze trajectory (Figure 1(a)) and is tasked to predict a likely continuation of the trajectory (Figure 1(b)). We expect video of correct action execution to represent normal user behavior, and hence to be characterized by predictable gaze patterns, while human behavior will be anomalous, and hence gaze will be unpredictable, when a mistake is made by the user. By comparing the predicted gaze with ground truth eye gaze trajectories estimated through a gaze tracker, we can determine whether a video segment is anomalous (Figure 1(c)).
Experiments on EPIC-Tent [5] show the effectiveness of the proposed method when compared with one-class anomaly detection methods [6, 7], and unsupervised mistake detection baselines.
In sum, the contributions of this work are as follows:
  • We investigate for the first time the problem of unsupervised mistake detection from egocentric video of human activity;
  • We propose a method to perform unsupervised mistake detection which leverages a gaze completion model to identify instances in which gaze patterns are unpredictable as a signal for human uncertainty;
  • We benchmark different approaches and baselines on the EPIC-Tent dataset, showing the effectiveness of the proposed method.
II Related Work
Our research is related to previous investigations on egocentric gaze estimation, use of gaze as an input signal in egocentric vision, mistake detection, and video anomaly detection.
II-A Egocentric gaze estimation
Previous works considered the problem of estimating where a person’s gaze falls in the scene from the observation of egocentric video. Since gaze characterizes human behavior [4], models able to estimate human gaze from visual observations prove useful in augmented reality and assistive technologies downstream applications even when a gaze tracker is not available [8, 9]. The seminal work of [10] investigated the simultaneous recognition of daily activities and prediction of gaze location patterns with a probabilistic generative model, showing the advantages of modeling gaze for the understanding of human behavior. The same work also introduced GTEA Gaze and GTEA Gaze+, two of the first datasets of daily activities including egocentric video and human gaze trajectories. Following up on this seminal work, other investigations proposed approaches for gaze prediction from egocentric video incorporating egocentric cues [11], modeling task-dependent attention transition [12], leveraging vanishing point, manipulation point, hand regions [13], and proposing specific architectural changes [14]. Other datasets to study egocentric gaze estimation and its applications in a variety of scenarios have been proposed through the years, leading to further development in gaze estimation approaches. These include EGTEA Gaze+ [15], a subset of Ego4D [16], EPIC-Tent [17], HoloAssist [2], and IndustReal[18]. Lai et al. recently proposed GLC [19], a transformer-based model integrating diverse gaze cues which achieves state-of-the-art performance on the EGTEA Gaze+ and Ego4D benchmarks.
In this work, we leverage the state-of-the-art GLC model [19] as an off-the-shelf approach to learn a model of egocentric human attention. We show that such a model, if trained with proper data sampling and optimization objectives, is instrumental for unsupervised mistake detection in egocentric activity. Additionally, we introduce the novel problem of gaze completion, where partial gaze trajectories are observed by the predicted model, and show that such an approach achieves superior performance.
II-B Use of gaze in egocentric vision
While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications. The seminal work of [8] exploited gaze for identifying task-relevant objects and their interaction modes from egocentric videos, achieving a 95% discovery rate across tasks. In [20], eye gaze movements are used to detect privacy-sensitive situations and preserve bystanders’ privacy. In [21], gaze is used to detect objects attended by the camera wearer. GazeGPT [22] utilizes eye tracking to assist large language models (LLMs) in classification tasks by discerning the user’s focal point within the world-facing camera view. GazeSAM [23] combines eye tracking with the Segment Anything Model for real-time object segmentation to allow radiologists to collect segmentation masks by looking at regions of interest during image diagnosis. GIMO [24] shows the advantages of incorporating eye gaze information for ego-centric human motion prediction. The work introduces a comprehensive dataset comprising body poses, scene scans, and ego-centric views with eye gaze, which are essential for accurately inferring human intent. TSM [25] is a hybrid text saliency model that merges cognitive reading models with human gaze supervision for Natural Language Processing (NLP). It showcases a robust correlation with human gaze across four datasets and suggests a joint modelling approach to incorporate TSM predictions into NLP tasks.
In this paper, we illustrate the critical role of gaze in mistake detection. Our method involves comparing gaze trajectories predicted from visual data with ground truth gaze estimate through a gaze tracker of the wearable device to identify mistakes when the model’s prediction deviates from the ground truth.
II-C Mistake Detection in Egocentric Videos
Mistakes naturally occur in human activities. The ability to automatically detect them from egocentric video can be beneficial for an AR assistant to offer support. A formal definition of this task was initially introduced in [3]. Identifying such mistakes usually entails modeling procedural knowledge [26], skill assessment [27], action segmentation [28] or detecting forgotten actions [29]. Notably, previous works tackled the task in a supervised fashion, training models to classify an action segment as “correct” or “mistake” on manually annotated instances [3, 2]. While this approach is feasible in closed-world scenario, it requires 1) a formal definition of what a mistake is, depending on the selected domain (e.g., kitchens vs the assembly line), 2) significant amounts of manually labeled data, which is expensive and requires expert knowledge. In this work, we tackle an unsupervised mistake detection task, in which models observe unlabeled video at training time and are tasked to output an anomaly score for given video segments at test time.
II-D Video Anomaly Detection
Our research also relates to the problem of Video Anomaly Detection (VAD), which involves recognizing abnormal or anomalous events within videos [6, 30]. A line of video anomaly methods are based on one-class classification, in which models are trained only on normal videos. These approaches usually aim to learn a dictionary of normal features, which can either be obtained based on hand-crafted features [31] or deep autoencoder models [32]. Weakly-supervised learning methods, like Multiple Instance Learning (MIL), leverage both normal and anomalous videos during training, utilizing video-level labels to predict anomalies at the segment level [33, 34, 35]. Notably, anomaly detection in egocentric vision remains underexplored [36]. Similar to video anomaly detection, we aim to detect mistakes by determining video segments which deviate from statistics observed at training time. Differently from previous works in video anomaly detection, we ground our predictions in an egocentric gaze estimation model, which acts as a proxy for modeling human behavior, hence effectively mistake prediction detection when anomalous behavior is observed.
Figure 2: We consider three approaches to compare the ground truth with respect to the predicted trajectories in order to determine a mistake. (a) Euclidean distance between the two trajectories. (b) Dynamic Time Warping (DTW). (c) Average value of the ground truth trajectory at the predicted heatmaps. .
III Proposed Approach
III-A Mistake Detection Problem Setup
The mistake detection task consists in highlighting those parts of the video in which the user is making a mistake during the execution of a given activity. In our setup, at each timestep $t$, a model $\Phi$ takes as input a video $V$ observed up to timestep $t$, $V_{1:t}$ and a 2D gaze trajectory $T_{1:t}$, where the $i-th$ element of the trajectory $T_{i}=(x_{i},y_{i})$ is a 2D gaze fixation, and have to return a score $s_{t}=\Phi(V_{1:t},T_{1:t})$ indicating whether a mistake is happening at the current time $t$. In this context, high $s_{t}$ scores indicate the occurrence of a mistake, while low $s_{t}$ scores indicate a correct action. We can hence see the mistake detection problem as a classification task, in which timesteps $t$ are classified as mistakes if $s_{t}>\phi$, where $\phi$ is a chosen threshold. We follow previous literature on anomaly detection [6, 7] and evaluate methods in a threshold-independent fashion by reporting the Receiver Operating Characteristics Area Under the Curve (ROC-AUC), where we consider “mistake” as the positive class [^1] In this context, a true positive is a mistake correctly classified as a mistake, a true negative is a correct execution correctly classified as a correct execution, a false positive is a correct execution wrongly classified as a mistake, and a false negative is a mistake wrongly classified as a correct execution.. For completeness, we also report the best $F_{1}$ score achieved considering the different threshold, as well as its related precision and recall values.
III-B Proposed Method
At each timestep $t$, we trim the input video $V_{1:t}$ and gaze trajectory $T_{1:t}$ obtained by the gaze tracker of the wearable device to the last observed $F$ frames, hence considering $V_{t-F:t}$ and $T_{t-F:t}$ as our inputs (Figure 1(a)). Our method relies on two main components: a gaze completion model (Figure 1(b)), and a scoring function (Figure 1(c)).
Figure 3: The model inputs consist of $t$ observed RGB frames denoted as $V_{t-F:t}$ and the 2D gaze trajectory represented by $T_{t-F:t}$. The first half of the trajectory, $T_{t-F:t-F/2}$, highlighted in orange, is utilized in one of the two gaze fusion strategies. Following a peak finding operation, the latter part of the trajectory, $T_{t-F/2:t}$, is predicted.
III-B1 Gaze Completion Module
[…]
Early and mid trajectory conditioning modules
Since GLC follows an encoder-decoder scheme, we experiment with two modules for injecting the input gaze trajectories into the model for the prediction of the output gaze trajectory.
The early fusion module adds the heatmaps in $Q_{1:F/2}$ to the first $F/2$ frames of the input video, $V_{t-F:t-F/2}$ as an additional channel. The values of this channel are set to zero for the remaining frames. This acts as a form of soft conditioning aiming to include information about the input gaze trajectory in the computation. Note that, in order to incorporate information into the input gaze, the model needs to learn how to compute suitable gaze representations from the additional input channel.
The mid fusion module processes the heatmaps $Q_{1:F/2}$ with a convolutional neural network encoder to produce a feature map, which is summed to the input of the decoder module of GLC. This module aims to encode gaze information in more direct way, with an additional convolutional neural network encoder components tasked to extract from the input gaze representations useful for GLC’s decoder.
Training Loss
Following [19], we consider gaze prediction as defining a probability distribution over the 2D image plane of each input frame. As in [19], we train the model by minimizing the sum of the Kullback–Leibler divergence between the predicted gaze maps $\hat{P}(i)$ and the ground truth ones $Q(i)$ at each frame $i$:
$L_{\text{KL}}(\hat{P}\parallel Q)=\sum_{i}\hat{P}(i)\log\left(\frac{\hat{P}(i)}{Q(i)}\right)$ (1)
III-B2 Scoring Function
Our method predicts the mistake confidence score $s_{t}$ by comparing the predicted trajectory $\hat{T}$ with the ground truth one $T_{t-F:t-F/2}$ which is obtained by the gaze tracker of the wearable device. We consider three different ways to compare the two trajectories: Euclidean distance, Dynamic Time Warping and heatmap.
Euclidean Distance
This method consists in using the Euclidean distance computed between corresponding points in each trajectory as the score $s_{t}$ (Fig. 2(a)):
$s_{t}=\sum_{i=0}^{F}\|T_{t-F+i}-\hat{T}_{i}\|$ (2)
where $\hat{T}_{i}=[\hat{x}_{t},\hat{y}_{t}]^{T}$ is predicted 2D gaze trajectory point at time step $t$.
Dynamic Time Warping
We then considered a more refined approach, utilizing Dynamic Time Warping (DTW) for comparison (Fig. 2(b)).
$DTW_{q}({T},{\hat{T}})=\min_{\pi\in\mathcal{A}({T},{\hat{T}})}\left(\sum_{(i,j)\in\pi}d(T_{i},\hat{T}_{j})^{q}\right)^{\frac{1}{q}}$ (3)
where $\pi$ is an alignment path, $\mathcal{A}({T},{\hat{T}})$ is the set of all admissible paths and $d$ is a distance measure.
Heatmap
We consider a probabilistic approach which evaluates the likelihood of a ground truth eye fixation $T_{i}$ obtained by the device under the predicted heatmap $\hat{P}$, which can be computed as:
$P(T_{i}|\hat{P})=\hat{P}(T_{i}^{x},T_{i}^{y})$ (4)
where $T_{i}^{x}$ and $T_{i}^{y}$ are the coordinates of the trajectory point $T_{i}=[T_{i}^{x},T_{i}^{y}]$. The score associated to the predicted trajectory $\hat{T}$ is computed as the sum of the likelihoods of each trajectory point, considering the predicted heatmap $\hat{P}$:
$s_{t}=\sum_{i=0}^{F}P(T_{t-F+i}|\hat{P})$ (5)
Entropy
This method consists of calculating the mean entropy of all predicted heatmaps for a given trajectory $\hat{T_{i}}$. The entropy $H$ of a single heatmap $\hat{P}$ is given by:
$H(\hat{P})=-\sum_{i}\hat{p}_{i}\log_{2}(\hat{p}_{i})$ (6)
where $p_{i}$ denotes the normalized probability of the $i$-th bin in the heatmap.
IV Benchmark Experiments
In this section we discuss the datasets used for the experiments, the implementation details, and the experimental results.
IV-A Datasets
IV-A1 EPIC-Tent
We base our main experiments on EPIC-Tent [17], a dataset comprising $7$ hours of egocentric video of $29$ subjects wearing two head-mounded cameras (GoPro and SMI eye tracker) while assembling a camping tent. The dataset includes egocentric video, gaze and labels indicating video segments in which users make mistakes. EPIC-Tent contains $151,821$ mistake frames and $384,558$ frames of correct execution, hence with a 28:72 ratio between correct and mistake frames. The unbalanced nature of the dataset arises from the natural rarity of mistakes and makes the problem particularly challenging to address. We randomly split videos in training, validation and test sets roughly following a 60:15:25 ratio.
IV-B Implementation Details
We follow [19] to set the hyperparameters of the GLC module. Differently from [19], we use a stride of $1$, rather than $2$ to train the model, which we found leads to marginal improvements in gaze and mistake prediction. To prevent overfitting, we set the weight decay to 0.07. We set the batch size to $4$ clips of $8$ frames.
IV-C Supervision Levels and Compared Approaches
We compared the proposed approach to methods belonging to two different classes involving different supervision levels: one-class classification methods, and unsupervised methods. Beyond compared methods, we also consider a random baseline in which a random score is assigned to each input clip.
IV-C1 One-Class Classification Methods
are trained only on videos of correct executions, following the standard anomaly detection setup [7, 6]. Note that, while these methods do not access any label, they still require the training set to include only correct executions, hence leading to a one-class setup rather than an unsupervised one. For this class, we compare our method with respect to TrajREC [6] and MoCoDAD [7], two popular approaches for video anomaly detection based on the processing of skeletal data. In the experiment, we replace skeletal data with gaze trajectories. We then consider an instantiation of the proposed approach in which the gaze completion model is trained only on correct executions, hence effectively replicating a one-class paradigm. Our intuition is that the gaze completion module will learn to better model normal behavior, hence leading to more variability in the predictions when a mistake is encountered. Finally, we compare with respect to a baseline which replaces the proposed gaze completion module with a simple gaze prediction component based on GLC [19]. Following the one-class setup, we train the GLC gaze predictor on correct executions and use the investigated scoring functions to predict scores by comparing the ground truth gaze trajectory obtained by the gaze tracker of the wearable device with respect to the predicted gaze.
IV-C2 Unsupervised Methods
assume as input an unlabeled set of clips including both mistakes and correct executions. We compare several variants of the proposed approach to unsupervised mistake detection based on gaze completion, exploring different scoring functions and trajectory conditioning modules, and compare with a gaze prediction baseline which replaces the gaze completion module with GLC and does not condition on partial trajectories.
IV-D Mistake detection on EPIC-Tent
TABLE I: Performance of various methods on EPIC-Tent with respect to the different supervision levels. Per-group best results are reported in bold.
Method Sup. Level Gaze at Test-Time Scoring Fun. AUC F1 Precision Recall
1 Random // 0.51 0.35 0.29 0.42
2 TrajREC [6] One-Class // 0.51 0.45 0.29 0.96
3 MoCoDAD [7] One-Class // 0.50 0.44 0.29 0.97
4 Gaze Prediction One-Class Entropy 0.55 0.41 0.28 0.78
5 Gaze Prediction One-Class Euclidean 0.60 0.41 0.29 0.69
6 Gaze Prediction One-Class DTW 0.61 0.43 0.30 0.74
7 Gaze Prediction One-Class Heatmap 0.66 0.46 0.37 0.62
8 Gaze Completion* One-Class Heatmap 0.66 0.34 0.42 0.29
[…]
11 Gaze Prediction Unsupervised Entropy 0.53 0.42 0.30 0.66
12 Gaze Prediction Unsupervised Euclidean 0.59 0.42 0.32 0.62
13 Gaze Prediction Unsupervised DTW 0.60 0.44 0.33 0.69
14 Gaze Prediction Unsupervised Heatmap 0.61 0.44 0.33 0.70
15 Gaze Completion* Unsupervised Heatmap 0.63 0.45 0.32 0.74
[…]
  • Early Fusion. Mid Fusion. Early and Mid score fusion.
Table I reports the results of the compared mistake detection approaches on the EPIC-Tent dataset, according to the different levels of supervision. For each method, we report the supervision level, whether the approach uses gaze at test time, the scoring function (where applicable) and the evaluation measures. One-class methods originally designed for anomaly detection from skeletal data (lines 2-3) do not improve over the random baseline with respect to the AUC score ($0.50-0.51$ vs $0.51$), and only marginally improve when considering the F1 score ($0.44-0.45$ vs $0.35$). These results suggests that only analyzing gaze trajectories, without relating them to the input video is insufficient for mistake detection. Using a gaze prediction model, paired with the entropy scoring function allows to achieve an AUC of $0.55$ and an F1 of $0.41$. It is worth noting that this approach does not require the availability of gaze at test time. Using different scoring functions which consider gaze as a test-time input allows to improve results in AUC to $0.6$, and $0.61$ when the Euclidean distance and DTW scoring functions are used. The improvement (from $0.55$ of line 4 to $0.61$ of line 6) highlights the usefulness of using gaze as an input signal at test time. Pairing gaze prediction with the heatmap scoring function further boosts the AUC to $0.66$, the F1 score to $0.46$, precision and recall to $0.37$ and $0.62$. These results show the feasibility of comparing predicted and ground truth gaze obtained by the gaze tracker at test time to detect mistakes. Adopting the proposed gaze completion approach allows to further boost results. In particular, this is observed when results of early fusion and mid fusion combined (line 10) [^3] We fuse the two techniques averaging the gaze prediction maps produced by the two models, allowing to achieve an AUC of $0.67$ and and F1 score of $0.47$, the largest score in this supervision level group. While the advantages of gaze completion over gaze prediction are limited in the one-class settings, gaze completion is much more effective than gaze prediction in the unsupervised setting. Indeed, in this case, using gaze completion with early fusion improves results over gaze prediction ($0.63$ of line 15 vs $0.61$ of line 14), and similarly with mid fusion ($0.62$ of line 16 vs $0.61$ of line 14). Our best model (line 17), combining early and mid fusion, achieves an AUC of $0.64$, an F1 score of $0.49$, a precision of $0.34$ and a recall of $0.88$ without seeing any mistake label. It has also a significant improvements over the random baseline (line 1: $0.51$ AUC, $0.35$ F1, $0.29$ precision, $0.42$ recall).
TABLE II: Comparison of GLC and GLC with Gaze Completion for Gaze Estimation on Epic-Tent.
Method F1 Recall Precision
GLC 0.37 0.65 0.26
Gaze Completion * 0.40 0.67 0.29
Gaze Completion 0.33 0.63 0.23
  • * Early Fusion. Mid Fusion
IV-E Gaze Estimation results
We have compared the performance of gaze prediction to those of gaze completion across the considered dataset. Table II reports results on the EPIC-Tent dataset. The Gaze Completion with Early Fusion (EF) method, exhibited superior performance, achieving the highest F1 score of 0.40, Recall of 0.67, and Precision of 0.29. This suggests that conditioning on partial trajectories makes the problem of making predictions more structured and hence easier to tackle.
Interestingly, while gaze completion with mid-fusion generally allows to improve over gaze prediction in the mistake detection problem, it does not improve gaze prediction results, suggesting that better mistake prediction is not necessarily achieved by methods able to make better gaze prediction.
Figure 4: Qualitative results of gaze completion approaches. Input gaze trajectories are highlighted in orange, predicted trajectories are in green for correct executions and in red for mistakes, ground truth trajectories to be predicted are in blue. Top: examples of correct executions. Our model predicts gaze trajectories (green) coherent with the ground truth (blue). Bottom: gaze is unpredictable when a mistake is made, leading to a large discrepancy between predicted (red) and ground truth (blue) trajectories.
IV-F Qualitative Results
Figure 4 reports qualitative results of the proposed gaze completion approach in the unsupervised setting. For correct executions (Figure 4-top), our model predicts a gaze trajectory (in green) similar to the ground truth one (in blue), taking as input the video and a partial trajectory (in orange). When a mistake is made by the camera wearer (Figure 4-bottom), our algorithm predicts gaze trajectories (in red) significantly different from the ground truth ones (in blue) as gaze becomes unpredictable.
V Conclusion
In this work, we have introduced an approach to detect mistakes in egocentric videos through unsupervised learning, exploiting the gaze signal as input. Grounded in the concept of ”gaze completion”, our approach predicts gaze trajectories based on observed video and partial gaze data. Detection of mistakes is enabled by comparing the predicted trajectory to the ground truth to identify moments in which gaze becomes unpredictable. Experimental validation, conducted on the EPIC-Tent dataset, demonstrates the efficacy of our proposed method, which outperforms existing one-class anomaly detection methods. The results of this study highlight the potential of utilizing gaze information, a readily available data source in many commercial devices, to enhance user experience. By proactively identifying and rectifying errors in task execution, this capability paves the way for more intuitive and interactive wearable technologies that adapt to the user’s needs in real-time.
Acknowledgements
Research at the University of Catania is supported by the project Future Artificial Intelligence Research (FAIR) – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006.

### v3

@@
-Eyes Wide Unshut: Unsupervised Mistake Detection in Egocentric Video by Detecting Unpredictable Gaze
+Eyes Wide Unshut: Unsupervised Mistake Detection in Egocentric Procedural Video by Detecting Unpredictable Gaze
 Abstract
-In this paper, we address the challenge of unsupervised mistake detection in egocentric video through the analysis of gaze signals, a critical component for advancing user assistance in smart glasses.
-Traditional supervised methods, reliant on manually labeled mistakes, suffer from domain-dependence and scalability issues.
-This research introduces an unsupervised method for detecting mistakes in videos of human activities, overcoming the challenges of domain-specific requirements and the necessity for annotated data.
-By analyzing unusual gaze patterns that signal user disorientation during tasks, we propose a gaze completion model that forecasts eye gaze trajectories from incomplete inputs.
-The difference between the anticipated and observed gaze paths acts as an indicator for identifying errors.
-Our method is validated on the EPIC-Tent dataset, showing its superiority compared to current one-class supervised and unsupervised techniques.
-Index Terms:
-Gaze, Mistake, Unsupervised, Behaviour Understanding
-I Introduction
+In this paper, we address the challenge of unsupervised mistake detection in egocentric procedural video through the analysis of gaze signals.
+Traditional supervised mistake detection methods rely on manually labeled mistakes, and hence suffer from domain-dependence and scalability issues.
+We introduce an unsupervised method for detecting mistakes in videos of human activities, overcoming the challenges of domain-specific requirements and the need for annotated data.
+We postulate that, when a subject is making a mistake in the execution of a procedure, their attention patterns will deviate from normality.
+We hence propose to detect mistakes by comparing gaze trajectories predicted from input video with ground truth gaze signals collected through a gaze tracker.
+Since predicting gaze in video is characterized by high uncertainty, we propose a novel gaze completion task, which aims to predict gaze from visual observations and partial gaze trajectories.
+We further contribute a gaze completion approach based on a Gaze-Frame Correlation module to explicitly model the correlation between gaze information and each local visual tokens.
+Inconsistencies between the predicted and observed gaze trajectories act as an indicator for identifying mistakes.
+Experiments on the EPIC-Tent, HoloAssist and IndustReal datasets showcase the effectiveness of the proposed approach as compared to unsupervised and one-class techniques.
+Our method is ranked first on the HoloAssist Mistake Detection challenge.
+Figure 1: We aim to detect mistakes in egocentric procedural video by analyzing video and gaze fixations acquired through a gaze tracker, assuming that gaze fixations are unpredictable in the case of mistakes.
+Left: ground truth gaze trajectory of a correct and wrong execution of a “place water container” action, together with gaze fixation maps averaged across many correct and wrong executions of the action.
+Note the higher variability exhibited by wrong executions.
+Right: The proposed unsupervised mistake detection method.
+(a) We assume as input a video with a partial gaze trajectory on the initial part of the video.
+(b) A gaze completion model predicts a gaze trajectory for the remaining part of the video, conditioned on the input video and the partial trajectory.
+(c) A mistake is detected if the predicted trajectory is significantly different from the observed one, which suggests an anomalous attention pattern.
+1 Introduction
 Smart glasses have recently gained significant popularity, with various existing products available on the market capable of providing assistance to the user through Augmented Reality.
@@
 Such a fully supervised approach has two main downsides: 1) It is domain-dependent, hence requiring an accurate characterization of what a mistake is, depending on the context.
-For instance, a mistake in a kitchen scenario is different from a mistake in the assembly line; 2) It requires a sufficient number of manually labeled mistake instances, which is time consuming, requires expert knowledge and sometimes it is difficult to observe and record.
+For instance, a mistake in a kitchen scenario is different from a mistake in the assembly line; 2) It requires to collect and label a sufficient number of mistake instances, which may be difficult to observe and record, involve time consuming procedures, and require expert knowledge.
 Ideally, a wearable assistant should be able to infer when the behavioral patterns of the user deviate from the norm in order to determine if they need assistance in a scenario-independent setting, i.e., without making specific assumptions on how a mistake is defined and without requiring costly labels.
-To overcome the limitations of previous approaches, we propose to detect mistakes in egocentric videos of human activity in an unsupervised way, by learning from unlabeled video.
-Figure 1: The proposed unsupervised mistake detection method.
-(a) We assume as input a video with a partial gaze trajectory on the initial part of the video.
-(b) A gaze completion model predicts a gaze trajectory for the remaining part of the video, conditioned on the input video and the partial trajectory.
-(c) A mistake is detected if the predicted trajectory is significantly different from the observed one, which suggests an anomalous attention pattern.
-We postulate that mistakes are characterized by abnormal behavior exhibited by the camera wearer, who may feel disoriented during the execution of the task at hand.
-As widely acknowledged in the literature, gaze patterns play an important role in the control of human activities [4], hence, we study how to analyze them as a means to perform mistake detection.
-We propose to train a model of human gaze behavior, in the form of a gaze predictor, and use it to predict likely eye gaze trajectories from video at inference time.
-Assuming that ground truth gaze trajectories are available as input, we frame gaze prediction as a novel “gaze completion” task in which a model takes as input a video and a partial gaze trajectory (Figure 1(a)) and is tasked to predict a likely continuation of the trajectory (Figure 1(b)).
-We expect video of correct action execution to represent normal user behavior, and hence to be characterized by predictable gaze patterns, while human behavior will be anomalous, and hence gaze will be unpredictable, when a mistake is made by the user.
-By comparing the predicted gaze with ground truth eye gaze trajectories estimated through a gaze tracker, we can determine whether a video segment is anomalous (Figure 1(c)).
-Experiments on EPIC-Tent [5] show the effectiveness of the proposed method when compared with one-class anomaly detection methods [6, 7], and unsupervised mistake detection baselines.
-In sum, the contributions of this work are as follows:
-  • We investigate for the first time the problem of unsupervised mistake detection from egocentric video of human activity;
-  • We propose a method to perform unsupervised mistake detection which leverages a gaze completion model to identify instances in which gaze patterns are unpredictable as a signal for human uncertainty;
-  • We benchmark different approaches and baselines on the EPIC-Tent dataset, showing the effectiveness of the proposed method.
-II Related Work
-Our research is related to previous investigations on egocentric gaze estimation, use of gaze as an input signal in egocentric vision, mistake detection, and video anomaly detection.
-II-A Egocentric gaze estimation
-Previous works considered the problem of estimating where a person’s gaze falls in the scene from the observation of egocentric video.
-Since gaze characterizes human behavior [4], models able to estimate human gaze from visual observations prove useful in augmented reality and assistive technologies downstream applications even when a gaze tracker is not available [8, 9].
-The seminal work of [10] investigated the simultaneous recognition of daily activities and prediction of gaze location patterns with a probabilistic generative model, showing the advantages of modeling gaze for the understanding of human behavior.
-The same work also introduced GTEA Gaze and GTEA Gaze+, two of the first datasets of daily activities including egocentric video and human gaze trajectories.
-Following up on this seminal work, other investigations proposed approaches for gaze prediction from egocentric video incorporating egocentric cues [11], modeling task-dependent attention transition [12], leveraging vanishing point, manipulation point, hand regions [13], and proposing specific architectural changes [14].
+To overcome limitations of previous approaches, we propose to detect mistakes in egocentric videos of human activity in an unsupervised way, learning from unlabeled video.
+We postulate that mistakes are characterized by abnormal attention patterns exhibited by the camera wearer, who may deviate from a standard procedure or feel disoriented during the execution of the task at hand.
+Think of a user trying to operate the coffee machine but forgetting to add water first.
+As the user pushes the brew button, they may notice that no coffee is being brewed and shift their attention back and forth from the cup to the water tank to the led light signaling the correct functioning of the machine, a significant deviation from a standard “button $\to$ cup $\to$ button” attention pattern.
+This observation is grounded in previous literature showing that gaze patterns play an important role in the control of human activities [17].
+We propose to learn a model of “normal” attention patterns, in the form of a gaze predictor capable of outputting likely eye gaze trajectories from a video at inference time.
+Since gaze prediction can be governed by high uncertainty, depending on the user’s goals, we propose a novel “gaze completion” task in which a model takes as input a video and a partial gaze trajectory (Figure 1(a)) and is tasked to predict a likely continuation of the trajectory (Figure 1(b)).
+Gaze completion is tackled with a novel approach based on a Gaze-Frame Correlation module to explicitly model the correlation between gaze information and each local visual token.
+We expect videos of correct action executions to represent normal user behavior, and hence to be characterized by predictable gaze patterns, while human behavior will deviate from normality, and hence gaze will be unpredictable when a mistake is made by the user.
+We hence predict mistakes by comparing the predicted gaze with the ground truth eye gaze trajectory obtained through a gaze tracker(Figure 1(c)).
+Experiments on EPIC-Tent [12], HoloAssist [36] and IndustReal[26], show the effectiveness of the proposed approach when compared to one-class anomaly detection methods [32, 7], and various unsupervised mistake detection baselines.
+In sum, the contributions of this work are as follows: 1) We investigate for the first time the problem of unsupervised mistake detection from egocentric video of human activity and provide an initial benchmark based on three different datasets, EPIC-Tent, HoloAssist, and IndustReal, and baseline results; 2) We define the novel “gaze completion” task consisting in predicting gaze trajectories from video and partial gaze inputs and introduce an approach to tackle this task based on a Gaze-Frame Correlation module; 3) We propose a method to perform unsupervised mistake detection which leverages a gaze completion model to identify instances in which gaze patterns are unpredictable as a signal for human uncertainty.
+Experiments show the effectiveness of the approach, while there is ample space for improvement in the newly proposed benchmark.
+We will publicly release the code and model checkpoints to support future research.
+2 Related Work
+Egocentric gaze estimation Literature on gaze estimation from egocentric video is rich [4, 18, 11, 34, 1, 19, 16].
+The seminal work of [4] investigated the simultaneous recognition of daily activities and prediction of gaze location patterns with a probabilistic generative model.
+Subsequent investigations proposed approaches for gaze prediction from egocentric video incorporating egocentric cues [18], modeling task-dependent attention transition [11], leveraging vanishing point, manipulation point, hand regions [34], and proposing specific architectural changes [1].
 Other datasets to study egocentric gaze estimation and its applications in a variety of scenarios have been proposed through the years, leading to further development in gaze estimation approaches.
@@
 recently proposed GLC [16], a transformer-based model integrating diverse gaze cues which achieves state-of-the-art performance on the EGTEA Gaze+ and Ego4D benchmarks.
-In this work, we leverage the state-of-the-art GLC model [19] as an off-the-shelf approach to learn a model of egocentric human attention.
-We show that such a model, if trained with proper data sampling and optimization objectives, is instrumental for unsupervised mistake detection in egocentric activity.
-Additionally, we introduce the novel problem of gaze completion, where partial gaze trajectories are observed by the predicted model, and show that such an approach achieves superior performance.
-II-B Use of gaze in egocentric vision
-While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications.
+We propose a novel gaze completion task and show its application to the problem of unsupervised mistake detection in egocentric video.
+We build on the state-of-the-art GLC model [16] to learn a model of egocentric human attention by tackling the proposed gaze completion task.
+Use of gaze in egocentric vision While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications.
 The seminal work of [2] exploited gaze for identifying task-relevant objects and their interaction modes from egocentric videos, achieving a 95% discovery rate across tasks.
@@
 GazeGPT [15] utilizes eye tracking to assist large language models (LLMs) in classification tasks by discerning the user’s focal point within the world-facing camera view.
-GazeSAM [23] combines eye tracking with the Segment Anything Model for real-time object segmentation to allow radiologists to collect segmentation masks by looking at regions of interest during image diagnosis.
-GIMO [24] shows the advantages of incorporating eye gaze information for ego-centric human motion prediction.
-The work introduces a comprehensive dataset comprising body poses, scene scans, and ego-centric views with eye gaze, which are essential for accurately inferring human intent.
-TSM [25] is a hybrid text saliency model that merges cognitive reading models with human gaze supervision for Natural Language Processing (NLP).
-It showcases a robust correlation with human gaze across four datasets and suggests a joint modelling approach to incorporate TSM predictions into NLP tasks.
-In this paper, we illustrate the critical role of gaze in mistake detection.
-Our method involves comparing gaze trajectories predicted from visual data with ground truth gaze estimate through a gaze tracker of the wearable device to identify mistakes when the model’s prediction deviates from the ground truth.
-II-C Mistake Detection in Egocentric Videos
-Mistakes naturally occur in human activities.
+GazeSAM [35] combines eye tracking with the Segment Anything Model [14] to allow radiologists to collect segmentation masks by looking at regions of interest during image diagnosis.
+GIMO [38] shows the advantages of incorporating eye gaze information for egocentric human motion prediction.
+TSM [29] is a hybrid text-saliency model merging cognitive reading models with human gaze for Natural Language Processing tasks.
+We show the effectiveness of gaze in mistake detection.
+Our method involves comparing gaze trajectories predicted from visual data with ground truth gaze estimated through a gaze tracker to identify mistakes when predictions deviates from the ground truth.
+Mistake Detection in Egocentric Videos Mistakes naturally occur in human activities.
 The ability to automatically detect them from egocentric video can be beneficial for an AR assistant to offer support.
 A formal definition of this task was initially introduced in [28].
-Identifying such mistakes usually entails modeling procedural knowledge [26], skill assessment [27], action segmentation [28] or detecting forgotten actions [29].
-Notably, previous works tackled the task in a supervised fashion, training models to classify an action segment as “correct” or “mistake” on manually annotated instances [3, 2].
-While this approach is feasible in closed-world scenario, it requires 1) a formal definition of what a mistake is, depending on the selected domain (e.g., kitchens vs the assembly line), 2) significant amounts of manually labeled data, which is expensive and requires expert knowledge.
-In this work, we tackle an unsupervised mistake detection task, in which models observe unlabeled video at training time and are tasked to output an anomaly score for given video segments at test time.
-II-D Video Anomaly Detection
-Our research also relates to the problem of Video Anomaly Detection (VAD), which involves recognizing abnormal or anomalous events within videos [6, 30].
+Identifying mistakes usually entails modeling procedural knowledge [3, 27], skill assessment [8], action segmentation [9] or detecting forgotten actions [30].
+Notably, previous works tackled the task in a supervised fashion, training models to classify an action segment as “correct” or “mistake” in manually annotated instances [28, 36].
+While this approach is feasible in a closed-world scenario, it requires 1) a definition of what a mistake is, depending on the domain (e.g., kitchens vs the assembly line), 2) significant amounts of manually labeled data, which is expensive and requires expert knowledge.
+In this work, we tackle an unsupervised mistake detection task, in which models observe unlabeled video at training time and are tasked to detect mistakes from video at test time.
+Our unsupervised scheme is possible through the analysis of gaze attention patterns, which provide a supervisory signal to create a joint video-gaze model of normal behavior.
+Video Anomaly Detection Our research also relates to the problem of Video Anomaly Detection (VAD), which involves recognizing abnormal or anomalous events within videos [32, 20].
 A line of video anomaly methods are based on one-class classification, in which models are trained only on normal videos.
-These approaches usually aim to learn a dictionary of normal features, which can either be obtained based on hand-crafted features [31] or deep autoencoder models [32].
+These approaches usually aim to learn a dictionary of normal features, which can either be obtained based on hand-crafted features [21] or deep models [37].
 Weakly-supervised learning methods, like Multiple Instance Learning (MIL), leverage both normal and anomalous videos during training, utilizing video-level labels to predict anomalies at the segment level [33, 22, 5].
-Notably, anomaly detection in egocentric vision remains underexplored [36].
+Notably, anomaly detection in egocentric vision remains under-explored [23].
 Similar to video anomaly detection, we aim to detect mistakes by determining video segments which deviate from statistics observed at training time.
-Differently from previous works in video anomaly detection, we ground our predictions in an egocentric gaze estimation model, which acts as a proxy for modeling human behavior, hence effectively mistake prediction detection when anomalous behavior is observed.
-Figure 2: We consider three approaches to compare the ground truth with respect to the predicted trajectories in order to determine a mistake.
-(a) Euclidean distance between the two trajectories.
-(b) Dynamic Time Warping (DTW).
-(c) Average value of the ground truth trajectory at the predicted heatmaps.
-.
-III Proposed Approach
-III-A Mistake Detection Problem Setup
+Differently from previous works in video anomaly detection, we ground our predictions in an egocentric gaze estimation model, which acts as a proxy for modeling normal human behavior, hence effectively achieving mistake prediction detection when anomalous behavior is observed.
+Moreover, we go beyond the one-class assumption and show that our method can also be used in an unsupervised setting in which unlabeled correct and mistake examples are included at training.
+3 Proposed Approach
+3.1 Mistake Detection Problem Setup
 The mistake detection task consists in highlighting those parts of the video in which the user is making a mistake during the execution of a given activity.
-In our setup, at each timestep $t$, a model $\Phi$ takes as input a video $V$ observed up to timestep $t$, $V_{1:t}$ and a 2D gaze trajectory $T_{1:t}$, where the $i-th$ element of the trajectory $T_{i}=(x_{i},y_{i})$ is a 2D gaze fixation, and have to return a score $s_{t}=\Phi(V_{1:t},T_{1:t})$ indicating whether a mistake is happening at the current time $t$.
+In our setup, at each time-step $t$, a model $\Phi$ takes as input a video $V$ observed up to time-step $t$, $V_{1:t}$ and a 2D gaze trajectory $T_{1:t}$, where the $i-th$ element of the trajectory $T_{i}=(x_{i},y_{i})$ is a 2D gaze fixation in frame $V_{i}$.
+Given this input, the model has to return a score $s_{t}=\Phi(V_{1:t},T_{1:t})$ indicating whether a mistake is happening at the current time $t$.
 In this context, high $s_{t}$ scores indicate the occurrence of a mistake, while low $s_{t}$ scores indicate a correct action.
@@
 We follow previous literature on anomaly detection [32, 7] and evaluate methods in a threshold-independent fashion by reporting the Receiver Operating Characteristics Area Under the Curve (ROC-AUC), where we consider “mistake” as the positive class [^1] In this context, a true positive is a mistake correctly classified as a mistake, a true negative is a correct execution correctly classified as a correct execution, a false positive is a correct execution wrongly classified as a mistake, and a false negative is a mistake wrongly classified as a correct execution..
-For completeness, we also report the best $F_{1}$ score achieved considering the different threshold, as well as its related precision and recall values.
-III-B Proposed Method
+For completeness, we also report the best $F_{1}$ score achieved considering the different thresholds, as well as its related precision and recall values.
+3.2 Proposed Method
 At each timestep $t$, we trim the input video $V_{1:t}$ and gaze trajectory $T_{1:t}$ obtained by the gaze tracker of the wearable device to the last observed $F$ frames, hence considering $V_{t-F:t}$ and $T_{t-F:t}$ as our inputs (Figure 1(a)).
 Our method relies on two main components: a gaze completion model (Figure 1(b)), and a scoring function (Figure 1(c)).
-Figure 3: The model inputs consist of $t$ observed RGB frames denoted as $V_{t-F:t}$ and the 2D gaze trajectory represented by $T_{t-F:t}$.
-The first half of the trajectory, $T_{t-F:t-F/2}$, highlighted in orange, is utilized in one of the two gaze fusion strategies.
-Following a peak finding operation, the latter part of the trajectory, $T_{t-F/2:t}$, is predicted.
-III-B1 Gaze Completion Module
+Figure 2: The model takes as input $F$ RGB frames $V_{t-F:t}$ and a partial 2D gaze trajectory $T_{t-F:t-F/2}$ on the first $F/2$ input frames, and outputs a predicted gaze heatmap $\hat{P}$ from the input video, conditioned on the trajectory.
+The input trajectory is encoded as a spatio-temporal heatmap $Q$.
+Trajectory and RGB inputs are fused using a combination of three strategies, channel fusion, which adds gaze heatmaps as a separate channel (1), correlation fusion, using a dedicated gaze-visual correlation module (2) and sum fusion, summing $Q$ to a mid-level representation of the input (3).
+We follow the design of [16] and process our inputs with a transformer encoder-decoder architecture which outputs a predicted gaze heatmap $\hat{P}$, supervised via a standard Kullback–Leibler divergence loss using the ground truth unobserved trajectory $T_{t-F/2:t}$.
+The output trajectory $\hat{T}_{t-F/2:t}$ is recovered from $\hat{P}$ using a peak finding operation.
+Gaze Completion Model Figure 2 illustrates the proposed gaze completion model.
+The model takes as input the video $V_{t-F:t}$ and the first half of the input gaze trajectory $T_{t-F:t-F/2}$ and predicts a gaze trajectory $\hat{T}_{t-F/2:t}$ aligned to the remaining part of the ground truth trajectory $T_{t-F/2:t}$.
+The goal of this model is to predict where the user is looking in the video, conditioned on an the initial trajectory.
+The conditioning allows to reduce the uncertainty on gaze predictions and give a prior into the intention of the user.
+We build on [16] and propose an encoder-decoder transformer-based architecture, experimenting with three approaches to condition gaze prediction on the input partial trajectory: channel fusion, correlation fusion, and sum fusion.
 […]
-Early and mid trajectory conditioning modules
-Since GLC follows an encoder-decoder scheme, we experiment with two modules for injecting the input gaze trajectories into the model for the prediction of the output gaze trajectory.
-The early fusion module adds the heatmaps in $Q_{1:F/2}$ to the first $F/2$ frames of the input video, $V_{t-F:t-F/2}$ as an additional channel.
+Channel Fusion The channel fusion module (Figure 2(1)) adds the heatmaps in $Q_{1:F/2}$ to the first $F/2$ frames of the input video, $V_{t-F:t-F/2}$ as an additional channel.
 The values of this channel are set to zero for the remaining frames.
-This acts as a form of soft conditioning aiming to include information about the input gaze trajectory in the computation.
-Note that, in order to incorporate information into the input gaze, the model needs to learn how to compute suitable gaze representations from the additional input channel.
-The mid fusion module processes the heatmaps $Q_{1:F/2}$ with a convolutional neural network encoder to produce a feature map, which is summed to the input of the decoder module of GLC.
-This module aims to encode gaze information in more direct way, with an additional convolutional neural network encoder components tasked to extract from the input gaze representations useful for GLC’s decoder.
-Training Loss
-Following [19], we consider gaze prediction as defining a probability distribution over the 2D image plane of each input frame.
-As in [19], we train the model by minimizing the sum of the Kullback–Leibler divergence between the predicted gaze maps $\hat{P}(i)$ and the ground truth ones $Q(i)$ at each frame $i$:
-$L_{\text{KL}}(\hat{P}\parallel Q)=\sum_{i}\hat{P}(i)\log\left(\frac{\hat{P}(i)}{Q(i)}\right)$ (1)
-III-B2 Scoring Function
-Our method predicts the mistake confidence score $s_{t}$ by comparing the predicted trajectory $\hat{T}$ with the ground truth one $T_{t-F:t-F/2}$ which is obtained by the gaze tracker of the wearable device.
-We consider three different ways to compare the two trajectories: Euclidean distance, Dynamic Time Warping and heatmap.
-Euclidean Distance
-This method consists in using the Euclidean distance computed between corresponding points in each trajectory as the score $s_{t}$ (Fig. 2(a)):
-$s_{t}=\sum_{i=0}^{F}\|T_{t-F+i}-\hat{T}_{i}\|$ (2)
-where $\hat{T}_{i}=[\hat{x}_{t},\hat{y}_{t}]^{T}$ is predicted 2D gaze trajectory point at time step $t$.
-Dynamic Time Warping
-We then considered a more refined approach, utilizing Dynamic Time Warping (DTW) for comparison (Fig. 2(b)).
-$DTW_{q}({T},{\hat{T}})=\min_{\pi\in\mathcal{A}({T},{\hat{T}})}\left(\sum_{(i,j)\in\pi}d(T_{i},\hat{T}_{j})^{q}\right)^{\frac{1}{q}}$ (3)
-where $\pi$ is an alignment path, $\mathcal{A}({T},{\hat{T}})$ is the set of all admissible paths and $d$ is a distance measure.
-Heatmap
-We consider a probabilistic approach which evaluates the likelihood of a ground truth eye fixation $T_{i}$ obtained by the device under the predicted heatmap $\hat{P}$, which can be computed as:
+This form of early fusion acts as a form of soft conditioning aiming to include information about the input gaze trajectory in the computation.
+Note that, in order to incorporate information into the input gaze, the model needs to learn how to compute suitable gaze representations from the additional input channel during training.
+Correlation Fusion This approach (Figure 2(2)) aims to fuse visual and gaze tokens inspired by global-local fusion originally proposed in [16].
+This is done in two stages.
+First, within the transformer encoder, where attention between all visual features and the gaze token is computed.
+In this stage, correlations between all visual tokens and across visual and gaze tokens are leveraged to obtain a strong representation.
+Second, within the dedicated gaze-visual correlation module.
+Here, attention is computed only between the gaze token and the visual token, which provides a much more focused mechanism to condition gaze prediction on the input trajectory.
+Note that this fusion mechanism operates both at the early (through the encoder) and mid (through the gaze-visual correlation module) levels.
+Sum Fusion The mid fusion module (Figure 2(3)) processes the input heatmaps $Q_{1:F/2}$ with a convolutional neural network encoder to produce a feature map which is summed to the input of the decoder transformer module.
+This module aims to inject gaze information at the mid level.
+Note that the CNN processing the input gaze heatmap is now tasked to extract from the input gaze representations useful for the decoder.
+Figure 3: We consider three approaches to compare the ground truth with respect to the predicted trajectories in order to determine a mistake.
+(a) Entropy.
+(b) Euclidean distance between the two trajectories.
+(c) Dynamic Time Warping (DTW).
+(d) Average value of the ground truth trajectory at the predicted heatmaps.
+Scoring Function Our method predicts the mistake confidence score $s_{t}$ by comparing the predicted trajectory $\hat{T}_{t-F/2:t}$ with the ground truth one $T_{t-F/2:t}$ which is obtained by the gaze tracker of the wearable device.
+We explore four different ways to compare the two trajectories: Euclidean distance, Dynamic Time Warping (DTW), Entropy, and Heatmap.
+These are illustrated in Figure 3 and described in the following.
+Euclidean Distance This method consists in using the Euclidean distance computed between corresponding points in each trajectory as the score $s_{t}$ (Fig. 3(a)):
+$s_{t}=\sum_{i=F/2+1}^{F}\|T_{t-F+i}-\hat{T}_{t-F+i}\|.$ (1)
+Dynamic Time Warping This scoring function uses Dynamic Time Warping (DTW)[25] to measure the distance between trajectories (Fig. 3(b)):
+$s_{t}=\min_{\pi\in\mathcal{A}({T},{\hat{T}})}\left(\sum_{(i,j)\in\pi}d(T_{i},\hat{T}_{j})^{q}\right)^{\frac{1}{q}}$ (2)
+where $\pi$ is an alignment path (a sequence of pairs of points belonging to the two trajectories minimizing the sum of distances), $\mathcal{A}({T},{\hat{T}})$ is the set of all admissible paths (all paths following an increasing temporal ordering) and $d$ is a distance measure (we use the Euclidean distance).
+The reader is referred to this link [^3] https://pypi.org/project/fastdtw/ for more details.
+Heatmap Differently from previous functions, this approach explicitly considers the probability values predicted by the model at each location (Fig.
+3(d)).
+Specifically, we evaluate the likelihood of a ground truth eye fixation $T_{i}$ obtained by the device under the predicted heatmap $\hat{P}$, which can be computed as:
 $p(T_{i}|\hat{P})=\hat{P}(T_{i}^{x},T_{i}^{y})$ (3)
@@
 The score associated to the predicted trajectory $\hat{T}$ is computed as the sum of the likelihoods of each trajectory point, considering the predicted heatmap $\hat{P}$:
-$s_{t}=\sum_{i=0}^{F}P(T_{t-F+i}|\hat{P})$ (5)
-Entropy
-This method consists of calculating the mean entropy of all predicted heatmaps for a given trajectory $\hat{T_{i}}$. The entropy $H$ of a single heatmap $\hat{P}$ is given by:
-$H(\hat{P})=-\sum_{i}\hat{p}_{i}\log_{2}(\hat{p}_{i})$ (6)
-where $p_{i}$ denotes the normalized probability of the $i$-th bin in the heatmap.
-IV Benchmark Experiments
-In this section we discuss the datasets used for the experiments, the implementation details, and the experimental results.
-IV-A Datasets
-IV-A1 EPIC-Tent
-We base our main experiments on EPIC-Tent [17], a dataset comprising $7$ hours of egocentric video of $29$ subjects wearing two head-mounded cameras (GoPro and SMI eye tracker) while assembling a camping tent.
+$s_{t}=\sum_{i=F/2+1}^{F}p(T_{t-F+i}|\hat{P}).$ (4)
+Entropy This is the only method which does not require ground truth gaze for computation.
+We consider this measure as a way to check whether mistakes are systematically characterized by uncertain gaze predictions (Fig.
+3(c)).
+In this case, the score $s_{t}$ is set as the mean entropy of all predicted heatmaps for a given trajectory $\hat{T_{i}}$.
+The entropy $H$ of a single heatmap $\hat{P}$ is given by:
+$s_{t}=-\frac{1}{F/2}\sum_{i=F/2+1}^{F}\sum_{x,y}\hat{P}_{t-F+i}(x,y)\log_{2}(\hat{P}_{t-F+i}(x,y))$ (5)
+where $P_{j}(x,y)$ is the value at coordinates $(x,y)$ of the heatmap $P_{j}$.
+4 Experiments and Results
+4.1 Datasets
+EPIC-Tent [13] includes $7$ hours of egocentric video of $29$ subjects wearing a head-mounded GoPro and an SMI eye tracker while assembling a camping tent.
 The dataset includes egocentric video, gaze and labels indicating video segments in which users make mistakes.
-EPIC-Tent contains $151,821$ mistake frames and $384,558$ frames of correct execution, hence with a 28:72 ratio between correct and mistake frames.
+EPIC-Tent contains $151,689$ mistake frames and $384,558$ frames of correct execution, hence with a 28:72 ratio between correct and mistake frames, which account to $86,099$, $27,613$, and $37,977$ mistake frames in training, validation, and test sets respectively.
 The unbalanced nature of the dataset arises from the natural rarity of mistakes and makes the problem particularly challenging to address.
 We randomly split videos in training, validation and test sets roughly following a 60:15:25 ratio.
-IV-B Implementation Details
-We follow [19] to set the hyperparameters of the GLC module.
+IndustReal [26] is designed for studying procedural tasks in industrial-like environments and consists of distinct training, validation, and test sets.
+The training set comprises $78,902$ frames, with $95.68\%$ frames labeled as correct and $4.32\%$ labeled as mistakes.
+The validation set includes $38,036$ frames, with $95.18\%$ correct and $4.82\%$ mistaken frames.
+The test set contains $90,105$ frames, with $92.53\%$ correct and $7.47\%$ mistaken frames.
+HoloAssist [36] focuses on a variety of scenarios in which users perform tasks with the assistance of an expert.
+The training set comprises $11,614,033$ frames, with $94\%$ frames labeled as correct and $6\%$ labeled as mistakes.
+The test set contains $1,699,562$ frames, with $95\%$ correct and $5\%$ mistaken frames.
+4.2 Implementation Details
+We base our model on the implementation provided in [16] and set hyperparameters accordingly.
+We will share the code to provide further details.
 Differently from [16], we use a stride of $1$, rather than $2$ to train the model, which we found leads to marginal improvements in gaze and mistake prediction.
@@
 We set the batch size to $4$ clips of $8$ frames.
-IV-C Supervision Levels and Compared Approaches
-We compared the proposed approach to methods belonging to two different classes involving different supervision levels: one-class classification methods, and unsupervised methods.
-Beyond compared methods, we also consider a random baseline in which a random score is assigned to each input clip.
-IV-C1 One-Class Classification Methods
-are trained only on videos of correct executions, following the standard anomaly detection setup [7, 6].
-Note that, while these methods do not access any label, they still require the training set to include only correct executions, hence leading to a one-class setup rather than an unsupervised one.
+Specifically, we process each video with non-overlapping windows of 8 consecutive frames.
+Each batch contains 4 windows (clips), for a total of 8*4=32 frames per batch.
+4.3 Supervision Levels and Compared Approaches
+We compare the proposed approach to methods belonging to two different supervision levels: one-class classification and unsupervised.
+Besides the compared methods, we also consider a random baseline in which a random score is assigned to each input clip.
+One-Class Classification Methods are trained only on videos of correct executions, following the standard setup of anomaly detection [7, 32].
+Note that, while these methods do not require any label during training, they still require the training set to include only correct executions, hence leading to a one-class setup rather than an unsupervised one.
 For this class, we compare our method with respect to TrajREC [32] and MoCoDAD [7], two popular approaches for video anomaly detection based on the processing of skeletal data.
-In the experiment, we replace skeletal data with gaze trajectories.
+In the experiments, we replace skeletal data with gaze trajectories.
 We then consider an instantiation of the proposed approach in which the gaze completion model is trained only on correct executions, hence effectively replicating a one-class paradigm.
-Our intuition is that the gaze completion module will learn to better model normal behavior, hence leading to more variability in the predictions when a mistake is encountered.
+Our intuition is that the gaze completion module will learn to model normal human behavior, hence leading to discrepant predictions when a mistake is encountered.
 Finally, we compare with respect to a baseline which replaces the proposed gaze completion module with a simple gaze prediction component based on GLC [16].
-Following the one-class setup, we train the GLC gaze predictor on correct executions and use the investigated scoring functions to predict scores by comparing the ground truth gaze trajectory obtained by the gaze tracker of the wearable device with respect to the predicted gaze.
-IV-C2 Unsupervised Methods
-assume as input an unlabeled set of clips including both mistakes and correct executions.
-We compare several variants of the proposed approach to unsupervised mistake detection based on gaze completion, exploring different scoring functions and trajectory conditioning modules, and compare with a gaze prediction baseline which replaces the gaze completion module with GLC and does not condition on partial trajectories.
-IV-D Mistake detection on EPIC-Tent
-TABLE I: Performance of various methods on EPIC-Tent with respect to the different supervision levels. Per-group best results are reported in bold.
-Method Sup. Level Gaze at Test-Time Scoring Fun. AUC F1 Precision Recall
-1 Random // 0.51 0.35 0.29 0.42
-2 TrajREC [6] One-Class // 0.51 0.45 0.29 0.96
-3 MoCoDAD [7] One-Class // 0.50 0.44 0.29 0.97
-4 Gaze Prediction One-Class Entropy 0.55 0.41 0.28 0.78
-5 Gaze Prediction One-Class Euclidean 0.60 0.41 0.29 0.69
-6 Gaze Prediction One-Class DTW 0.61 0.43 0.30 0.74
-7 Gaze Prediction One-Class Heatmap 0.66 0.46 0.37 0.62
-8 Gaze Completion* One-Class Heatmap 0.66 0.34 0.42 0.29
-[…]
-[…]
-11 Gaze Prediction Unsupervised Entropy 0.53 0.42 0.30 0.66
-12 Gaze Prediction Unsupervised Euclidean 0.59 0.42 0.32 0.62
-13 Gaze Prediction Unsupervised DTW 0.60 0.44 0.33 0.69
-14 Gaze Prediction Unsupervised Heatmap 0.61 0.44 0.33 0.70
-15 Gaze Completion* Unsupervised Heatmap 0.63 0.45 0.32 0.74
-[…]
-[…]
-  • Early Fusion. Mid Fusion. Early and Mid score fusion.
-Table I reports the results of the compared mistake detection approaches on the EPIC-Tent dataset, according to the different levels of supervision.
-For each method, we report the supervision level, whether the approach uses gaze at test time, the scoring function (where applicable) and the evaluation measures.
-One-class methods originally designed for anomaly detection from skeletal data (lines 2-3) do not improve over the random baseline with respect to the AUC score ($0.50-0.51$ vs $0.51$), and only marginally improve when considering the F1 score ($0.44-0.45$ vs $0.35$).
-These results suggests that only analyzing gaze trajectories, without relating them to the input video is insufficient for mistake detection.
-Using a gaze prediction model, paired with the entropy scoring function allows to achieve an AUC of $0.55$ and an F1 of $0.41$.
-It is worth noting that this approach does not require the availability of gaze at test time.
-Using different scoring functions which consider gaze as a test-time input allows to improve results in AUC to $0.6$, and $0.61$ when the Euclidean distance and DTW scoring functions are used.
-The improvement (from $0.55$ of line 4 to $0.61$ of line 6) highlights the usefulness of using gaze as an input signal at test time.
-Pairing gaze prediction with the heatmap scoring function further boosts the AUC to $0.66$, the F1 score to $0.46$, precision and recall to $0.37$ and $0.62$.
-These results show the feasibility of comparing predicted and ground truth gaze obtained by the gaze tracker at test time to detect mistakes.
-Adopting the proposed gaze completion approach allows to further boost results.
-In particular, this is observed when results of early fusion and mid fusion combined (line 10) [^3] We fuse the two techniques averaging the gaze prediction maps produced by the two models, allowing to achieve an AUC of $0.67$ and and F1 score of $0.47$, the largest score in this supervision level group.
-While the advantages of gaze completion over gaze prediction are limited in the one-class settings, gaze completion is much more effective than gaze prediction in the unsupervised setting.
-Indeed, in this case, using gaze completion with early fusion improves results over gaze prediction ($0.63$ of line 15 vs $0.61$ of line 14), and similarly with mid fusion ($0.62$ of line 16 vs $0.61$ of line 14).
-Our best model (line 17), combining early and mid fusion, achieves an AUC of $0.64$, an F1 score of $0.49$, a precision of $0.34$ and a recall of $0.88$ without seeing any mistake label.
-It has also a significant improvements over the random baseline (line 1: $0.51$ AUC, $0.35$ F1, $0.29$ precision, $0.42$ recall).
-TABLE II: Comparison of GLC and GLC with Gaze Completion for Gaze Estimation on Epic-Tent.
-Method F1 Recall Precision
-GLC 0.37 0.65 0.26
-Gaze Completion * 0.40 0.67 0.29
-Gaze Completion 0.33 0.63 0.23
-  • * Early Fusion. Mid Fusion
-IV-E Gaze Estimation results
-We have compared the performance of gaze prediction to those of gaze completion across the considered dataset.
-Table II reports results on the EPIC-Tent dataset.
-The Gaze Completion with Early Fusion (EF) method, exhibited superior performance, achieving the highest F1 score of 0.40, Recall of 0.67, and Precision of 0.29.
-This suggests that conditioning on partial trajectories makes the problem of making predictions more structured and hence easier to tackle.
-Interestingly, while gaze completion with mid-fusion generally allows to improve over gaze prediction in the mistake detection problem, it does not improve gaze prediction results, suggesting that better mistake prediction is not necessarily achieved by methods able to make better gaze prediction.
-Figure 4: Qualitative results of gaze completion approaches.
-Input gaze trajectories are highlighted in orange, predicted trajectories are in green for correct executions and in red for mistakes, ground truth trajectories to be predicted are in blue.
-Top: examples of correct executions.
-Our model predicts gaze trajectories (green) coherent with the ground truth (blue).
-Bottom: gaze is unpredictable when a mistake is made, leading to a large discrepancy between predicted (red) and ground truth (blue) trajectories.
-IV-F Qualitative Results
+Following the one-class setup, we train the GLC gaze predictor on correct executions and use the considered scoring functions to predict scores by comparing the ground truth gaze trajectory obtained by the gaze tracker of the wearable device with respect to the predicted gaze.
+Unsupervised Methods area trained on an unlabeled set of clips including both mistakes and correct executions.
+We evalaute different variants of the proposed approach to unsupervised mistake detection based on gaze completion, exploring different scoring functions and trajectory conditioning modules, and compare with a gaze prediction baseline based on GLC [16].
+Table 1: Ablation study results for various scoring functions and fusion strategies on EPIC-Tent in the unsupervised setting.
+Best results per-block are underlined, while best global results are in bold.
+CH: channel fusion, CORR: correlation fusion, S: sum fusion.
+Scoring Fusion F1 Precision Recall AUC
+1 Random // 0.36 0.29 0.42 0.51
+2 Entropy // 0.41 0.27 0.62 0.51
+3 Euclidean // 0.42 0.29 0.60 0.55
+4 DTW // 0.44 0.31 0.68 0.56
+5 Heatmap // 0.45 0.32 0.70 0.57
+6 Heatmap CH 0.45 0.32 0.74 0.63
+7 Heatmap CORR 0.5 0.36 0.82 0.65
+8 Heatmap S 0.48 0.35 0.82 0.62
+9 Heatmap CH + CORR 0.51 0.36 0.85 0.69
+10 Heatmap CH + S 0.49 0.35 0.82 0.65
+11 Heatmap CORR + S 0.49 0.35 0.81 0.64
+12 Heatmap CH + CORR + S 0.47 0.34 0.8 0.64
+4.4 Ablation Study
+Table 1 reports an ablation study evaluating the considered scoring functions and gaze-video fusion strategies in the unsupervised mistake detection on EPIC-Tent.
+In rows 2-5, we compare all scoring functions when all fusion strategies are turned off and the model is not conditioned on previous trajectories.
+The entropy scoring function achieved an AUC of 0.51 and an F1 score of 0.41 (row 1), only marginally above the random baseline, leading to the conclusion that high entropy in the prediction only partially correlates with the presence of a mistake.
+Notably, this approach does not require gaze data at test time.
+Employing alternative scoring functions that consider gaze as a test-time input improved the results, yielding an AUC of 0.55 and 0.56 when using Euclidean distance and DTW scoring functions, respectively (rows 2-3).
+The heatmap-based scoring function produced the best results in this comparison, with an AUC of 0.57 and an F1 score of 0.45, significantly above random level (compare with the F1 score of $0.36$ in row 1).
+The advantage of the random baseline is likely due to the better exploitation of the probability values computed by the gaze prediction model, as compared to other scoring functions.
+Hence, we adopted the heatmap-based scoring as our primary method in following comparisons.
+Rows 6-8 compare approaches using one of the three fusion strategies.
+While all fusion strategies improve results (compare row 6-8 with 5), the proposed correlation strategy (CORR) systematically outperforms channel and sum fusion, obtaining an AUC score of $0.65$ and an F1 score of $0.5$ and doubling the recall of the random baseline ($0.82$ vs $0.42$) with a better precision ($0.36$ vs $0.29$).
+Rows 9-12 finally explore combinations of fusion strategies.
+The best results are obtained by combining our proposed Correlation fusion with Channel fusion, achieving an AUC of 0.69 and an F1 score of 0.51 (+0.12 and +0.06 compared to the standard heatmap method - row 5).
+This configuration is the one referred to as “ours” in future comparisons, unless stated otherwise.
+Interestingly, adding the Sum fusion strategy leads to worse results.
+We speculate that this is due to the challenges associated with sum fusion, which requires an expressive mechanism to map trajectory information to the mid-level representations of the transformer encoder.
+Moreover, we can see the correlation strategy as both an early fusion method (at the token embedding layer) and as a mid fusion method (at the correlation level), which likely allows correlation fusion to capture similar information as sum fusion.
+Table 2: Comparison of GLC and the proposed Gaze Completion approach for Gaze Estimation on EPIC-Tent.
+Method Fusion F1 Recall Precision
+GLC // 0.37 0.65 0.26
+Gaze Completion CH 0.40 0.67 0.29
+Gaze Completion CH + CORR 0.43 0.70 0.31
+Table 2 compares the performance of GLC [16] used for standard gaze prediction with respect to different versions of our gaze completion approach on the EPIC-Tent dataset.
+Just using channel fusion brings a performance boost, achieving an F1-score of 0.40, recall of 0.67, and precision of 0.29, while combining channel and correlation fusion brings best results with an F1-score of 0.43, recall of 0.70, and precision of 0.31.
+These results suggest that conditioning on partial trajectories makes the problem of gaze prediction less uncertain and the proposed approach can leverage the informative content provided by the input trajectory surpassing the performance of a standard gaze prediction method.
+Moreover, performance in Table 2 correlates with performance in Table 1, suggesting that good gaze prediction is essential to mistake detection performance.
+Correlation between self-reported uncertainty and mistakes: Our main assumption is that gaze patterns can signal the presence of a mistake due to the user feeling disoriented and exhibiting abnormal attention patterns.
+We conducted a logistic regression analysis to explore the relationship between self-reported uncertainty levels (normalized in a $1-5$ range) included in the EPIC-Tent dataset [13] and the occurrence of labeled mistakes.
+Our analysis revealed that an increase in self-reported uncertainty by 1 unit is associated with a doubling the odds of the occurrence of a mistake  [^4] We obtain a coefficient of 0.68 with a $p<0.001$..
+Figure 5 illustrates the distributions of self-reported uncertainty values with respect to the occurrence of mistakes.
+Table 3: Mistake detection result on EPIC-Tent. Best results are in bold.
+Method Sup. Level F1 Precision Recall AUC
+Random // 0.36 0.29 0.42 0.51
+TrajREC [32] One-Class 0.45 0.29 0.96 0.51
+MoCoDAD [7] One-Class 0.44 0.29 0.97 0.50
+GLC [16] One-Class 0.46 0.37 0.62 0.66
+Ours One-Class 0.52 0.37 0.85 0.69
+GLC [16] Unsupervised 0.44 0.33 0.70 0.61
+Ours Unsupervised 0.51 0.36 0.85 0.69
+4.5 Comparison with Competitors
+Table 3 compares the proposed approach with competitors on the EPIC-Tent, according to the two considered levels of supervision.
+One-class methods such as TrajREC and MoCoDAD, originally designed for anomaly detection from skeletal data, show minor improvements over the random baseline in terms of AUC score ($0.50-0.51$ vs $0.51$) and only small improvement in F1 score ($0.44-0.45$ vs $0.35$).
+High recall values, paired with low precision, suggest that such methods tend to classify most clips as mistakes.
+These results indicate that solely analyzing gaze trajectories without considering the input video is insufficient for effective mistake detection.
+The state-of-the-art gaze prediction approach, GLC [16], shows improved results compared to the one-class methods, highlighting the advantages of incorporating gaze analysis for mistake detection.
+The proposed method brings best results yielding an AUC of 0.69, F1-score of 0.52, precision of 0.37, and recall of 0.85, improvements of $+3/6\%$ with respect to the best approach GLC and $+10/20\%$ with respect to the random baseline.
+While the benefits of gaze completion over gaze prediction are small but systematic in one-class settings, gaze completion proves significantly more effective in unsupervised scenarios.
+Indeed, in this case, our method brings significant improvements over both the random baseline (e.g., $0.69$ vs $0.51$ AUC score and $0.51$ vs $0.36$ F1 score) and the GLC competitor with $+8\%$ in AUC score, $+7\%$ in F1 score, $+3\%$ in precision and $+15\%$ in recall.
+Table 4: Mistake detection result on HoloAssist.
+Method Sup. Level F1 Precision Recall AUC
+Random // 0.04 0.02 0.4 0.5
+TrajREC [32] One-Class 0.09 0.04 0.96 0.5
+MoCoDAD [7] One-Class 0.11 0.06 0.94 0.51
+GLC [16] One-Class 0.19 0.11 0.56 0.60
+Ours One-Class 0.22 0.14 0.59 0.61
+GLC [16] Unsupervised 0.10 0.06 0.34 0.54
+Ours Unsupervised 0.18 0.12 0.40 0.59
+Table 5: Mistake detection result on IndustReal.
+Method Sup. Level F1 Recall Precision AUC
+Random // 0.12 0.06 0.62 0.51
+TrajREC [32] One-Class 0.18 0.10 0.90 0.56
+MoCoDAD [7] One-Class 0.19 0.11 0.91 0.57
+GLC [16] One-Class 0.21 0.15 0.33 0.60
+Ours One-Class 0.24 0.18 0.35 0.63
+GLC [16] Unsupervised 0.21 0.15 0.33 0.58
+Ours Unsupervised 0.21 0.16 0.33 0.62
+Figure 4: Uncertainty boxplot
+EPIC-Tent
+IndustReal
+HoloAssist
+Figure 5: Failure cases by action classes
+Table 4 compares the proposed method with competitors on HoloAssist.
+The HoloAssist dataset presents a more varied and expansive context compared to EPIC-Tent, including different sorts of activties, which makes it significantly more challenging to perform mistake detection.
+This is evidenced by the fact that the Random baseline achieves an F1-score of only $0.04$.
+One-class baselines like TrajREC and MoCoDAD bring improvement over the baseline but suffer from a tendency to overly classify clips as mistakes (Recall $0.96$ and $0.94$).
+GLC brings significant improvements in AUC and F1 scores, with slighlty more balanced precision and recall values.
+Best results are obtained by our approach, which achieves an AUC of 0.61 and an F1-score of 0.22 in one-class settings.
+Performances are slightly lower in the unsupervised case with an AUC score of $0.59$ and an F1 score of $0.18$, which still exhibit significant improvements over GLC, with gains of $+5\%$ and $+8\%$ respectively.
+Figure 6: Qualitative results of the proposed approach.
+Input gaze trajectories are highlighted in orange, predicted trajectories are in green for correct executions and in red for mistakes, ground truth trajectories are in blue.
+Table 5 reports results on the IndustReal dataset.
+Results confirm previous observations with TrajREC and MoCoDAD bringing minor improvements over the random baseline, and our method outperforming all competitors and improving over GLC by systematic margings both the one-class case ($+3\%$ in AUC and $+4\%$ in F1) as well as in the unsupervised one ($+4\%$ in AUC score and comparable F1 performance).
+4.6 Qualitative Results and Failure Cases
 Figure 6 reports qualitative results of the proposed gaze completion approach in the unsupervised setting.
 For correct executions (Figure 6-top), our model predicts a gaze trajectory (in green) similar to the ground truth one (in blue), taking as input the video and a partial trajectory (in orange).
-When a mistake is made by the camera wearer (Figure 4-bottom), our algorithm predicts gaze trajectories (in red) significantly different from the ground truth ones (in blue) as gaze becomes unpredictable.
-V Conclusion
-In this work, we have introduced an approach to detect mistakes in egocentric videos through unsupervised learning, exploiting the gaze signal as input.
-Grounded in the concept of ”gaze completion”, our approach predicts gaze trajectories based on observed video and partial gaze data.
-Detection of mistakes is enabled by comparing the predicted trajectory to the ground truth to identify moments in which gaze becomes unpredictable.
-Experimental validation, conducted on the EPIC-Tent dataset, demonstrates the efficacy of our proposed method, which outperforms existing one-class anomaly detection methods.
-The results of this study highlight the potential of utilizing gaze information, a readily available data source in many commercial devices, to enhance user experience.
-By proactively identifying and rectifying errors in task execution, this capability paves the way for more intuitive and interactive wearable technologies that adapt to the user’s needs in real-time.
-Acknowledgements
-Research at the University of Catania is supported by the project Future Artificial Intelligence Research (FAIR) – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006.
+When a mistake is made by the camera wearer (Figure 6-bottom), our algorithm predicts gaze trajectories (in red) significantly different from the ground truth ones (in blue).
+Figure 5 reports failure cases by action classes in EPIC-Tent (a), IndustReal (b) and HoloAssist (c).
+Most failures are with actions requiring fine eye-hand coordination (insert support, align) or actions not involving interactions (instruction, inspect), few failures are observed in simple actions (pick up/open/place, take, place).
+5 Conclusion
+In this work, we introduced an approach for detecting mistakes in egocentric videos in an unsupervised way, leveraging gaze signals as input.
+We proposed a novel gaze completion task, where we predict gaze trajectories based on observed video and partial gaze data and an approach to tackle this task.
+Detection of mistakes is performed by comparing predicted trajectories against ground truth, identifying instances where gaze behavior becomes unpredictable as potential mistakes.
+Experimental validation on the EPIC-Tent, HoloAssist, and IndustReal datasets demonstrates the efficacy of our method, surpassing traditional one-class anomaly detection techniques and other unsupervised mistake detection methods.
+Our method is ranked first on the HoloAssist Mistake Detection challenge.
+Limitations.
+Without strong assumptions on the semantics of the task, the proposed method is able to detect certain classes of mistakes, mainly related to fine hand-object coordination, while it may miss other classes of mistakes such as those related to key-step ordering for which procedure-aware methods may be beneficial.